A tank used for filling helium balloons has a volume of 0.600 m³ and contains 3.00 mol of helium gas at 30.0ºC. Assume the helium behaves like an ideal gas


1
Expert's answer
2021-05-19T11:05:54-0400

Solution.

"V=0.600m^3;"

"\nu=3.00mol;"

"T=303K;"

"pV=\nu RT\implies p=\dfrac{\nu RT}{V};"

"p=\dfrac{3.00mol\sdot8.31J/(molK)\sdot303K}{0.600m^3}=12589.65Pa;"


Answer:"p=12589.65Pa."